# From Calculation to Adjudication: Examining LLM judges on Mathematical Reasoning Tasks

### v1

@@ 9 Conclusion
We conducted a thorough analysis of LLM judges on mathematical reasoning tasks. We include a detailed judgment performance evaluation of eight models on three datasets. We find that larger models are generally better than smaller models and that judges succeed in detecting the more capable model. Our analysis reveals a strong correlation between judgment performance and task performance of the models providing candidate answers which shows that judges tend to choose larger or better models. We hypothesize that LLM judges incorporate writing style into their judgments instead of purely analyzing the reasoning. We provide two experiments to provide evidence for this hypothesis.
Finally, we want to emphasize the importance of impartiality and fairness in the role of LLM judges, similar to human judges in the real world. Our research introduces methods to quantify biases in favor of larger or better models, thereby offering a means to measure the reduction of such biases.
Limitations
Our analysis is primarily focused on mathematical reasoning datasets, which allows us to explore judgments through the lens of correctness within specific subsets. While this approach provides valuable insights, it limits the generalizability of our findings to other tasks or domains. Based on the fact that the investigated datasets are complex, in the sense that they need multi-step reasoning to be solved, and based on the fact that there is no thorough investigation of LLM judges on mathematical reasoning datasets yet, we think this work is a valuable contribution.
In our experiments, we focus on testing a single, specific prompt. It is common knowledge that LLMs are highly sensitive to variations in prompt phrasing, which can substantially influence their performance. Nevertheless, it is impossible for us to meet the computational demands necessary to run our experiments with multiple prompts.
In this study, we intentionally concentrate on open-weight models, motivated by our strong belief in the principles of open science. Open-weight models offer transparency and reproducibility, which are critical for advancing scientific understanding. However, we note that it is also interesting to study closed models to understand potential differences. Still, we are committed to research on open-weight models because we believe it benefits the community more.
10 Acknowledgements
This research was funded by the WWTF through the project ”Knowledge-infused Deep Learning for Natural Language Processing” (WWTF Vienna Research Group VRG19-008) and through the project "Transparent and Explainable Models" (WWTF ICT19-041). MA is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of BERD@NFDI - grant number 460037581. Further, we thank Jan Philip Wahle and Pedro Henrique Luz de Araujo for fruitful discussions and their constructive feedback.
@@ A.3 Prompts
Initial Prompt { "role": "user", "content": "You are a reasoning assistant. Always answer exactly in the same format. Use ’####’ to separate the final answer (without additional comments) from the reasoning. {{shot 1 question}}" }, { "role": "assistant", "content": "{{shot 1 question}}" } , { "role": "assistant", "content": "{{shot 4 answer}}" }, { "role": "user", "content": "{{question}}" } }
Figure 7: Prompt used to generate initial solutions for all datasets. It includes few-shots and the question of the current sample.
@@
Judge Prompt Question: {{question}} Answer A: {{answer A}} ————– Answer B: {{answer B}} ————– Compare both answers in detail and choose the answer which correctly answers the question. Conclude with a JSON in Markdown format indicating your choice between answer A or B: “‘json { "answer": "B" or "A" } “‘
Figure 8: Prompt used for judgements. The full text above is wrapped in the user role, as all models support this role. No additional system message is used.
@@
We used two different prompts within this project. The prompt shown in Figure 7 is used for the candidate solutions for all datasets. Examples of the few-shots are in Table 6. The prompt for the judges is given in Figure 8. Note that we run experiments for both orders of the answers of the models $M_{A}$ and $M_{B}$.
Judge model $M_{A}$ model $M_{B}$ Label Prompt Response Correct
Qwen 2 72B Qwen 2 72B Llama 3 8B 20.00 Question: Charlie wants to sell beeswax candles. For every pound of beeswax, he can make 10 tapered candles. One pound of beeswax and the wicks cost \$10.00 in supplies. If he sells each candle for \$2.00 each, what is his net profit if he makes and sells 20 candles? Answer A: He needs 2 pounds of beeswax to make 20 candles because 20 / 10 = <<20/10=2>>2 It will cost \$20.00 in supplies for the beeswax and wicks because 2 x 10 = <<2*10=20>>20 His total revenue from selling the candles will be \$40.00 because 20 x 2 = <<20*2=40>>40 His net profit will be \$20.00 because 40 - 20 = <<40-$20 ###$ ————– Answer B: It takes 1 pound of beeswax and the wicks to make 10 candles, so it takes 1 pound of beeswax and the wicks to make 20 candles. It costs 10 dollars for 1 pound of beeswax and the wicks, so it costs 10 dollars for 20 candles. Charlie sells each candle for 2 dollars, so he makes 20 * 2 = 40 dollars by selling the 20 candles. So, his net profit is 40 - 10 = 30 dollars. #### 30 ————– Compare both answers in detail and choose the answer which correctly answers the question. Conclude with a JSON in Markdown format indicating your choice between answer A or B: “‘json { "answer": "B" or "A" } “‘ Answer A and Answer B differ in their calculation of the cost of supplies for making 20 candles. Answer A correctly calculates that Charlie needs 2 pounds of beeswax to make 20 candles and that the cost for the beeswax and wicks for 2 pounds is \$20.00. Answer B incorrectly assumes that the cost for making 20 candles is the same as for 10 candles, which is \$10.00. Therefore, Answer A correctly calculates the net profit as \$20.00, while Answer B incorrectly calculates it as \$30.00. The correct answer is A. “‘json { "answer": "A" } “‘ True
Llama 3 70B Qwen 2 72B Llama 3 8B D Question: There is well of depth 30m and frog is at bottom of the well. He jumps 3m up one day and falls back 2m down the same day. How many days will it take for the frog to come out of the well? Options: A)25 days B)26 days C)27 days D)28 days E)29 days Answer A: The frog is initially at the bottom of the well, which is 30m deep. On the first day, the frog jumps 3m up, so it is 33m from the bottom. Then, it falls back 2m, so it is 31m from the bottom. Each day, the frog jumps 3m up and falls 2m down, making a net gain of 1m. To get out of the 30m deep well, it would take 30 days for the frog to jump and fall 30 times, covering a total distance of 30m. #### A ————– Answer B: If he jumps 3m up and slips 2m down in a day, he is 1m up in a day in 27 days he would have reached 27m On the 28th day, he jumps up 3m and come out of the well. Answer: D #### D ————– Compare both answers in detail and choose the answer which correctly answers the question. Conclude with a JSON in Markdown format indicating your choice between answer A or B: “‘json { "answer": "B" or "A" } “‘ Let’s break down both answers: **Answer A:** The explanation is correct that the frog makes a net gain of 1m each day. However, the conclusion that it would take 30 days to get out of the 30m deep well is incorrect. The frog would reach the top of the well when it has covered a total distance of 30m, but it would not take 30 days. **Answer B:** This explanation is correct. The frog makes a net gain of 1m each day, so in 27 days, it would have covered 27m. On the 28th day, it would jump up 3m and come out of the well. Therefore, the correct answer is **D) 28 days**. Here is the conclusion in JSON format: “‘json { "answer": "B" } “‘ True
Table 8: Example of ground truth answers used for few-shot prompting.
@@ A.4 Infrastructure
The experiments were run on NVIDIA A100 and NVIDIA H100. The judgments used in Section 4 and Section 5 took around 5 days equivalents on 4 A100 40GB. Using 2 H100 90GB and 4 A100 40GB it took around 2.5 days. For the perturbation experiments in Section 7, it took around the same amount of time.
@@ B.1 Task Performance
AQUA-RAT GSM8K MATH
Qwen 2 72B 76.38 92.04 51.19
Llama 3 70B 73.62 91.05 34.37
Yi 1.5 34B 64.96 78.47 27.04
Mixtral 8x7B 47.24 61.18 13.79
Llama 3 8B 51.18 73.01 15.04
Gemma 1.1 7B 42.91 50.72 12.60
Mistral 7B v0.3 38.19 42.76 6.13
Mistral 7B v0.1 21.65 26.08 3.10
Table 9: Task performance of all models using the prompt in Figure 7.
@@ B.2 Judging performance per model pair
(a) Qwen 2 70B
(b) LLama3 70B
(c) LLama 3 8B
(d) Gemma 1.1 7B
(e) Yi 1.5 34B
(f) Mixtral 8x7B
(g) Mistral 7B v0.3
(h) Mistral 7B v0.1
Figure 9: Evaluation of final task performance $P(\Delta_{J}=T|A,B)$ averaged over all datasets for model pairs $(M_{A},M_{B})$ for the judges (a) - (h).
We conduct experiments with all eight models serving as judges. We present the performance metrics of all judges across various model comparisons in Figure 9. As shown in Table 1, only the large models consistently produce judgments that deviate consistently from random chance. These results in Figure 9 support the superior performance of larger models.
Appendix C Additional subset experiments
This section provides additional information for chapter 5.
@@ C.1 Example Subset performance
To better understand the correlation observed in Figure 5, we provide examples of these subsets, which can be seen in Table 10. These examples include the following details: the judge, the compared models, the dataset, the performance of each model on a dataset (denoted by $P(A|D)$), the judgment performance on the subset (denoted by $P(\Delta_{J}=T|A=X,B=Y,D)$), and the relative size of the subset (denoted by $P(A=X,B=Y|D)$). We provide the five subsets with the highest performance, the five subsets with the five lowest performance, and five random subsets.
Judge model A model B dataset X Y $P(A|D)$ $P(B|D)$ $P(\Delta_{J}=T|A=X,B=Y,D)$ $P(A=X,B=Y|D)$
Qwen 2 72B Qwen 2 72B Mistral 7B v0.1 MATH True False 51.2 3.1 99.1 50.2
Qwen 2 72B Yi 1.5 34B Mistral 7B v0.1 MATH True False 27.0 3.1 98.4 27.9
Qwen 2 72B Llama 3 8B Mistral 7B v0.1 MATH True False 15.0 3.1 98.3 16.7
Qwen 2 72B Llama 3 70B Mistral 7B v0.1 MATH True False 34.4 3.1 98.3 35.4
Qwen 2 72B Mixtral 8x7B Mistral 7B v0.1 MATH True False 13.8 3.1 98.2 15.4
Mixtral 8x7B Mixtral 8x7B Mixtral 8x7B GSM8K False True 61.2 61.2 65.8 14.7
Yi 1.5 34B Llama 3 8B Gemma 1.1 7B GSM8K False True 73.0 50.7 64.9 8.0
Yi 1.5 34B Llama 3 70B Mistral 7B v0.1 AQUA-RAT True False 73.6 21.7 89.1 60.8
Qwen 2 72B Gemma 1.1 7B Gemma 1.1 7B GSM8K False True 50.7 50.7 90.0 12.9
Yi 1.5 34B Yi 1.5 34B Mistral 7B v0.1 GSM8K False True 78.5 26.1 52.5 2.6
Qwen 2 72B Llama 3 70B Mistral 7B v0.1 MATH False True 34.4 3.1 13.2 2.2
Qwen 2 72B Yi 1.5 34B Mistral 7B v0.1 AQUA-RAT False True 65.0 21.7 10.0 4.7
Yi 1.5 34B Qwen 2 72B Mistral 7B v0.1 MATH False True 51.2 3.1 6.5 1.4
Llama 3 70B Qwen 2 72B Mistral 7B v0.1 MATH False True 51.2 3.1 6.2 1.3
Qwen 2 72B Qwen 2 72B Mistral 7B v0.1 MATH False True 51.2 3.1 6.1 1.4
Table 10: Examples of comparisons; and performance; problem:
C.2 Performance by agreement
(a) Qwen 2 72B: Using all comparisons.
(b) Comparison with a correct and incorrect answer
(c) Llama 3 70B: Using all comparisons.
(d) Comparison with a correct and incorrect answer
(e) Yi 1.5 34B: Using all comparisons.
(f) Comparison with a correct and incorrect answer
(g) Mixtral 8x7B: Using all comparisons.
(h) Comparison with a correct and incorrect answer
Figure 10: Judge performance by agreement bucket, e.g. bucket 3 (X-axis) means that all eight models gave together three different answers.
We provide an extension of the results in the main paper. We put all samples into bins of how many different answers were given by the eight models. For the sake of brevity, we average across all judges. In Figure 10 we show the same graphs for each judge individually. Again, we give the performance on the overall dataset and on the subset where there exists exactly one correct and one incorrect solution. All interpretations and conclusions overlap with the main part of the paper.
Appendix D Prediction of Individual Judgements
Random Forest.
We utilize TF-IDF representations with English stop word removal for the Random Forest model. We further remove all digits from the text. We set the maximum number of features to 10,000 and consider n-grams ranging from unigrams to bigrams (N-gram range: 1-2). The model uses 500 trees (estimators). For training, we use the Scikit-learn Pedregosa et al. (2011) library. The running time was negligible.
RoBERTa.
For the RoBERTa model Liu et al. (2020), we use a batch size of 64 and a learning rate of 2e-5. The weight decay is set to 1e-3, and the model is trained for 8 epochs. The final model is selected based on the best validation performance. The model is trained using the HuggingFace Transformers library Wolf et al. (2020). The total running time was about twelve hours on a single H100 90GB.
Appendix E Statistical Methodology
@@ E.3 Multiple Testing
Appendix F Peturbation of Results
In Section 7 we performed two experiments, where we 1) swap the results of the individual answers and 2) mask all numbers using an ’X’ to understand whether the model only focuses on the writing style. You can find examples, including the prompt and the response of the judge for experiment 1) in Table 11 and examples for experiment 2) in Table 12
Judge model $M_{A}$ model $M_{B}$ Label Prompt Response Correct
Qwen 2 72B Qwen 2 72B Llama 3 8B 163.00 Question: Candice put 80 post-it notes in her purse before she headed out to her job at the coffee shop. On her way, she stopped off at the store and purchased a package of Post-it notes; At work, she placed a single Post-it note on each of 220 different cups of coffee. If she had 23 post-it notes remaining overall, how many Post-it notes were in the package that she purchased? Answer A: Candice had 80 post-it notes initially. She used 80 + 220 = <<80+220=300>>300 of them to note the cups of coffee. She had 23 post-it notes remaining overall. So, she used 300 - 23 = <<300-23=277>>277 of the 300 post-it notes initially. The remaining post-it notes are the difference between the total and the initial count, which is 300 - 80 = <<300-80=220>>220. She must have purchased 220 - 23 = <<220-23=163>>197 post-it notes. #### 163 ————– Answer B: Candice started with 80 Post-it notes and placed one on each of the 220 cups of coffee, so she used 220 notes. If she had 23 notes remaining, then she must have had 220 + 23 = 243 notes in total. Since she originally had 80 notes, the package she purchased must have contained 243 - 80 = 197 notes. Therefore, the package contained 197 Post-it notes. #### 197 ————– Compare both answers in detail and choose the answer which correctly answers the question. Conclude with a JSON in Markdown format indicating your choice between answer A or B: “‘json { "answer": "B" or "A" } “‘ Answer A is incorrect because it incorrectly calculates the number of post-it notes Candice used and purchased. The correct answer is B because it accurately calculates the total number of post-it notes Candice had after using them and then subtracts the initial number of post-it notes she had to find out how many were in the package she purchased. “‘json { "answer": "B" } “‘ True
Llama 3 70B Qwen 2 72B Llama 3 8B D Question: In a chocolate store, all chocolates are either vanilla or cocoa flavored only. 10% of the chocolates are cocoa flavored, 90% of the rest are squashed. What percentage of the chocolates are both vanilla flavored and not squashed? Options: A)1% B)2% C)5% D)9% E)10% Answer A: If 10% of the chocolates are cocoa flavored, then the remaining 90% are vanilla flavored. 90% of the vanilla chocolates are squashed. This means that 10% of the vanilla chocolates are not squashed. The question asks for the percentage of chocolates that are both vanilla flavored and not squashed. Since 10% of the vanilla chocolates are not squashed, the answer is 10%. #### D ————– Answer B: 90% of the chocolates are not cocoa flavored (they are vanilla) and 10% of these (90% *10% = 9%) are not squashed. So 9% of the chocolates are both vanilla flavored and not squashed. The answer is E. #### E ————– Compare both answers in detail and choose the answer which correctly answers the question. Conclude with a JSON in Markdown format indicating your choice between answer A or B: “‘json { "answer": "B" or "A" } “‘ Let’s break down both answers: **Answer A:** If 10% of the chocolates are cocoa flavored, then the remaining 90% are vanilla flavored. 90% of the vanilla chocolates are squashed. This means that 10% of the vanilla chocolates are not squashed. The question asks for the percentage of chocolates that are both vanilla flavored and not squashed. Since 10% of the vanilla chocolates are not squashed, the answer is 10%. **Issue with Answer A:** The calculation "10% of the vanilla chocolates are not squashed" is incorrect. 90% of the vanilla chocolates are squashed, which means 10% of 90% (not 10% of the total) are not squashed. **Answer B:** 90% of the chocolates are not cocoa flavored (they are vanilla) and 10% of these (90% * 10% = 9%) are not squashed. So 9% of the chocolates are both vanilla flavored and not squashed. **Correctness of Answer B:** This answer correctly calculates the percentage of chocolates that are both vanilla flavored and not squashed. Therefore, the correct answer is **B**. Here is the conclusion in JSON format: “‘json { "answer": "B" } “‘ True
Table 11: Examples of comparisons where the final results were swapped.
Judge model $M_{A}$ model $M_{B}$ Label Prompt Response Correct
[…]
Llama 3 70B Qwen 2 72B Llama 3 8B D Question: There is well of depth 30m and frog is at bottom of the well. He jumps 3m up one day and falls back 2m down the same day. How many days will it take for the frog to come out of the well? Options: A)25 days B)26 days C)27 days D)28 days E)29 days Answer A: The frog is initially at the bottom of the well, which is Xm deep. On the first day, the frog jumps Xm up, so it is Xm from the bottom. Then, it falls back Xm, so it is Xm from the bottom. Each day, the frog jumps Xm up and falls Xm down, making a net gain of Xm. To get out of the Xm deep well, it would take X days for the frog to jump and fall X times, covering a total distance of Xm. #### A ————– Answer B: If he jumps Xm up and slips Xm down in a day, he is Xm up in a day in X days he would have reached Xm On the Xth day, he jumps up Xm and come out of the well. Answer: D #### D ————– Compare both answers in detail and choose the answer which correctly answers the question. Only analyze the reasoning! Therefore we exchanged all numbers with ’X’ so you can focus on the reasoning. Conclude with a JSON in Markdown format indicating your choice between answer A or B: “‘json { "answer": "B" or "A" } “‘ Let’s analyze the reasoning of both answers: **Answer A:** The answer assumes that the frog makes a net gain of Xm (3m - 2m = 1m) each day. However, it doesn’t consider the fact that the frog needs to reach the top of the well to come out. The answer simply states that it would take X days to cover a total distance of Xm, without considering the remaining distance to the top of the well. **Answer B:** The answer correctly assumes that the frog makes a net gain of Xm (3m - 2m = 1m) each day. It then concludes that in X days, the frog would have reached Xm, and on the Xth day, it would jump up Xm and come out of the well. This answer considers the remaining distance to the top of the well and correctly concludes that the frog would come out on the Xth day. Based on the analysis, I choose: “‘json { "answer": "B" } “‘ True
Table 12: Example of comparisons where numbers were masked using an ’X’.

### v2

@@ 8 Conclusion
We conduct a thorough analysis of LLM judges on mathematical reasoning tasks. We evaluate the judgment performance of eight models of different sizes on three datasets. We find that larger judge models generally outperform smaller judge models and that judges can reliably detect whether both answers are correct. Our analysis reveals a strong correlation between judgments and task performance, indicating that judges tend to choose models of higher quality even if their answers are incorrect. We hypothesize that LLM judges partially base their decisions on linguistic cues in contrast to the reasoning within the answers. We support this hypothesis with our experiments showing that 70% of the judges’ decisions can be predicted using simple linguistic features such as N-grams of part-of-speech tags. Lastly, our analysis finds that LLM judges reliably detect LLMs of higher task performance but are not reliably useable to improve task performance. Our results show that LLM judges contain biases and suggest that practitioners should not blindly trust LLM judges. We advise practitioners to carefully decide whether LLM judges should be used in their particular application.
With this work, we set the stage for further research to investigate how to understand, use, and improve LLM judges.
9 Limitations
Our analysis is primarily focused on mathematical reasoning datasets, which allows us to explore judgments through the lens of verifiability, i.e., problems that have a definitely correct answer. While this approach provides valuable insights, it limits the generalizability of our findings to other tasks or domains. Nevertheless, we want to emphasize the importance of the class of verifiable tasks. For instance, there is currently a focus on training so-called large reasoning models, which demonstrate significant progress in solving complex problems such as coding or maths. It is a possibility that an increased capability of LLMs on verifiable tasks fuels scientific progress.
In our experiments, we focus on testing a single, specific prompt. It is common knowledge that LLMs are highly sensitive to variations in prompt phrasing, which can substantially influence their performance. However, the resources available to us do not allow us to meet the computational demands necessary to run our experiments with multiple prompts. Further, our impression is that it is a custom approach to conduct LLM studies using single prompts, as they are typically indicative of behavior. Therefore we decided to run our analysis on full datasets with a single prompt instead of using subsets of datasets with variations of the prompt with mostly the same content.
@@ A.3 Prompts
User You are a reasoning assistant. Always answer exactly in the same format. Use ’####’ to separate the final answer (without additional comments) from the reasoning. << Few-Shot Question 1 >>
Assistant << Few-Shot Answer 1 >>
User << Few-Shot Question 4 >>
Assistant << Few-Shot Answer 4 >>
User << Sample Question >>
Figure 7: The prompt to solve tasks. Few-shots and actual questions are filled in within “<<” and “>>” symbols.
@@
User Question: << question >> Answer A: << answer A >> ————– Answer B: << answer B >> ————– Compare both answers in detail and decide whether both answers are correct, both answers are incorrect or whether answer 1 or answer 2 is correct. Conclude with a JSON in Markdown format indicating your choice between "answer_1", "answer_2", "both_correct" or "both_incorrect": “‘json { "answer": "…" } “‘
Figure 8: Judge Prompt. Candidate answers are filled in within “<<” and “>>” symbols.
@@
We used two different prompts within this project. In general, we designed the prompts to be minimial, by assigning a minimal personality, a quick task description, and description of the output format. The prompt shown in Figure 7 is used for the candidate solution generation for all datasets. Examples of the few-shots are in Table 5. The prompt for the judges is given in Figure 8. Note that we run experiments for both orders of the answers of the models $A$ and $B$.
@@ A.4 Infrastructure
The experiments were run on NVIDIA A100 and NVIDIA H100. The judgments used in Section 4 took around 3 day equivalents on 4 A100 40GB. Using 2 H100 90GB and 4 A100 40 GB it took less than 2 days.
@@ B.1 Task Performance
GSM8K AQUA-RAT MATH
Llama 3.1 70B 93.25 78.57 47.11
Qwen 2.5 72B 95.07 83.73 73.86
Qwen 2.5 14B 93.48 82.54 64.47
Gemma 2 27B 85.97 67.46 38.80
Qwen 2.5 7B 88.10 75.40 60.31
Gemma 2 9B 80.52 61.51 31.31
Llama 3.1 8B 72.40 61.51 20.69
Gemma 2 2B 37.53 26.98 7.15
Table 7: Task performance of the investigated models.
@@ B.2 Judging performance per model pair
(a) LLama 3.1 70B
(b) Qwen 2.5 72B
(c) Qwen 2.5 14B
(d) Gemma 2 27B
(e) Qwen 2.5 7B
(f) Gemma 2 9B
(g) LLama 3.1 8B
(h) Gemma 2 2B
Figure 9: Performance $S^{J}_{A,B}$ of LLM judges on model pairs, averaged across datasets.
We conduct experiments with all eight models serving as judges. We present the performance metrics of all judges across all model pairs in Figure 9.
Appendix C Examples
@@ C.1 Example Subset Performance
Judge model A model B dataset $S_{A}$ $S_{B}$ $S^{J}_{A,B}$ No. Samples
Llama 3.1 70B Qwen 2.5 14B Gemma 2 2B MATH 64.50 7.10 94.60 1655
Llama 3.1 70B Qwen 2.5 72B Gemma 2 2B AQUA-RAT 83.70 27.00 94.50 309
Llama 3.1 70B Qwen 2.5 7B Gemma 2 2B MATH 60.30 7.10 94.00 1520
Llama 3.1 70B Qwen 2.5 72B Qwen 2.5 7B AQUA-RAT 83.70 75.40 62.50 64
Llama 3.1 70B Qwen 2.5 7B Qwen 2.5 14B AQUA-RAT 75.40 82.50 42.30 26
Llama 3.1 70B Qwen 2.5 72B Qwen 2.5 72B MATH 73.90 73.90 49.50 206
Llama 3.1 70B Gemma 2 27B Qwen 2.5 14B AQUA-RAT 67.50 82.50 15.00 20
Llama 3.1 70B Gemma 2 2B Qwen 2.5 14B GSM8K 37.50 93.50 15.00 20
Llama 3.1 70B Gemma 2 2B Llama 3.1 70B MATH 7.10 47.10 14.50 62
Table 8: Examples of judgement performances on subsets where model $A$ is correct and model $B$ is incorrect.
To better understand the correlation observed in Figure 4, we provide examples of these subsets, which can be seen in Table 8. These examples include the following details: the judge, the compared models, the dataset, the performance of the correct model on the dataset (denoted by $S_{A}$), the performance of the incorrect model on the dataset $S_{B}$, the judgment performance on the subset (denoted by $S^{J}_{A,B}$), and the size of the subset. We provide the three subsets with the highest performance, the three subsets with the lowest performance, and three random subsets where Llama 3.1 70B is the judge.
Appendix D Statistical Methodology
@@ D.3 Multiple Testing
Appendix E Sample-level Analysis
We utilize Scikit-learn Pedregosa et al. (2011) library to train and evaluate the Logistic Regression and Random Forest Model. We use the standard settings for the Logistic Regression model. We use the Random Forests model with 1500 estimators, and standard settings apart from that.
During preprocessing, we use simple word splittling by spaces. We employ the english stop word removal integreated into Scikit-learn. We set the maximum number of features to 5,000, for the N-Gram of part-of-speech tags, we set the N-gram range from 5-grams up to 13-grams, following settings of Shaib et al. (2024). For training, we use the Scikit-learn Pedregosa et al. (2011) library. The running time was negligible.
In Table 9
Model Both correct A correct B correct Both incorrect
Llama 3.1 70B 51.8 18.1 21.7 8.4
Qwen 2.5 72B 54.9 19.6 19.8 5.6
Qwen 2.5 14B 50.2 20.0 23.0 6.9
Gemma 2 27B 52.6 15.9 15.4 16.0
Table 9: Percentage of predictions individual models made.